### v1

@@ 5 Summary and outlook
Using an expansion in Laguerre polynomials Bassetto:1987fq and following the calculations for gluon jets in ref. Dokshitzer:1993dc , we obtain the KNO scaling functions for quark and gluon jets by incorporating energy conservation into the Double Logarithmic Approximation (DLA) of the generating function evolution equation. The resulting modified DLA (MDLA) expressions deviate markedly from the DLA results, supporting earlier conclusions that the latter fail to provide a reliable description of experimental data Malaza:1984vv ; Dokshitzer:1993dc ; Duan:2025ngi ; Dokshitzer:2025owq ; Dokshitzer:2025fky , although recent DLA-based studies have successfully explained the observed KNO scaling in DIS Liu:2022bru ; Liu:2023eve . For gluon jets, our results are consistent with ref. Dokshitzer:1993dc , while for both quark and gluon jets they exhibit quantitative differences compared with the recently suggested QCD-inspired expressions in ref. Dokshitzer:2025fky .
@@
The MDLA KNO scaling function for gluon jets depends solely on the anomalous dimension $\gamma_{0}$, while the quark KNO scaling function additionally depends on the ratio of mean multiplicities of quark and gluon jets in the asymptotic high-virtuality limit, $c_{q}$. Inspired by recent studies in refs. Dokshitzer:2025owq ; Dokshitzer:2025fky , we treat $c_{q}$ as a fitting parameter in our analysis of jets in $pp$ collisions, rather than using the DLA value of $C_{F}/C_{A}$. We find that by fixing $c_{q}$ and $\gamma_{0}$, the MDLA results can very well reproduce the approximate KNO scaling functions for both quark and gluon jets in the two leading jets of $pp$ collisions at $\sqrt{s}=13$ TeV, as extracted using PYTHIA for a broad jet $p_{T}$ range of 0.10.12.52.5 TeV in ref. Duan:2025ngi . Furthermore, we observe that combining the MDLA scaling functions with the leading-order cross section and the N3LO mean multiplicities in refs. Dremin:1999ji ; Capella:1999ms provides a very good description of the inclusive charged-particle multiplicity distributions measured by ATLAS ATLAS:2019rqw .
@@
Describing the inclusive multiplicity distributions requires additional theoretical input, such as the fraction of jets initiated by each parton type and mean multiplicities. This extra dependence may introduce uncertainties when evaluating the validity of KNO scaling. To address this and as a consistency check, we utilize jet substructure techniques to distinguish quark and gluon jets. Specifically, we apply the topic modeling method Metodiev:2018ftz ; Komiske:2018vkc to two jet samples provided by ATLAS ATLAS:2019rqw to extract multiplicity distributions for jet topics, serving as an approximation for quark and gluon jets. We find that these results are consistent with our MDLA results, although the experimental uncertainties remain significant.
@@
Several aspects of our study could be further improved. First, a complete calculation of the KNO scaling functions for quark and gluon jets beyond the DLA approximation is still lacking. This is important because, although MDLA only incorporates energy conservation compared to DLA, it already introduces significant modifications. A key issue to address beyond DLA is how to perform a direct calculation using the same jet definition (e.g., the anti-$k_{t}$ algorithm) employed in experiments. For inclusive multiplicity distributions, higher-order dijet cross sections are available Currie:2017eqf . In addition, the discrepancies between theoretical predictions and the measured mean multiplicities need to be resolved to achieve better agreement. One promising approach for high-precision calculations involves infrared-safe multiplicity definitions Medves:2022ccw ; Medves:2022uii , although multiplicity distributions within this approach are yet to be studied. Notably, KNO scaling functions are infrared safe within both DLA and MDLA, and parton-level results have been used for comparisons with experimental data in this work (see also refs. Dokshitzer:2025owq ; Dokshitzer:2025fky ). At the same time, non-perturbative effects are observed in PYTHIA simulations Vertesi:2020utz ; Duan:2025ngi , highlighting the importance of clarifying the role of hadronization effects Azimov:1984np ; Kang:2023zdx . Another promising avenue is to explore how studies of multiplicity distributions in quark and gluon jets can benefit from complementary methods of quark–gluon discrimination Gallicchio:2011xq ; Gallicchio:2012ez ; Larkoski:2014pca ; Gras:2017jty ; Metodiev:2017vrx ; Dreyer:2021hhr . On the experimental side, more precise measurements are needed, and the $n_{ch}$ binning should be refined.
@@ Acknowledgements.
X.D. thanks Dr. Wen-Ya Wu for valuable discussions regarding the treatment of experimental data. This work is supported by the European Research Council under project ERC-2018-ADG-835105 YoctoLHC; by Maria de Maeztu excellence unit grant CEX2023-001318-M and projects PID2020-119632GB-I00 and PID2023-152762NB-I00 funded by MICIU/AEI/10.13039/501100011033; and by ERDF/EU. It has received funding from Xunta de Galicia (CIGUS Network of Research Centres). X.D. is supported by the China Scholarship Council under Grant No. 202306100165. L.C. is supported by the Marie Skłodowska-Curie Actions Postdoctoral Fellowships under Grant No. 101210595. X.D. and G.M. are supported by the National Natural Science Foundation of China under Grants No.12147101, No. 12325507, the National Key Research and Development Program of China under Grant No. 2022YFA1604900, and the Guangdong Major Project of Basic and Applied Basic Research under Grant No. 2020B0301030008. B.W. acknowledges the support of the Ramón y Cajal program with the Grant No. RYC2021-032271-I and the support of Xunta de Galicia under the ED431F 2023/10 project.
@@ Appendix A Mean multiplicities at N3LO
To clarify the role of the parameters appearing in the N3LO results Dremin:1999ji ; Capella:1999ms , we provide a summary in this appendix. The mean multiplicity for gluon jets takes the form
@@
$\delta_{q}(y)=\delta_{g}(y)+\frac{C}{\sqrt{y}}r_{1}+\frac{C^{2}}{y}(r_{2}+\frac{r_{1}^{2}}{2}).$ (52)
The coefficients $a_{i},r_{i}$ depend on the number of flavors $n_{f}$ and are tabulated in refs. Capella:1999ms . In our analysis, these coefficients are evaluated using $n_{f}=5$, with $a_{1}=0.314,a_{2}=-0.301,a_{3}=0.112,r_{1}=0.198,r_{2}=0.510$ and $r_{3}=-0.041$.

### v2

@@ 5 Summary and outlook
Using an expansion in Laguerre polynomials Bassetto (1988) and following the calculations for gluon jets in ref. Dokshitzer (1993), we obtain the KNO scaling functions for quark and gluon jets by incorporating energy conservation into the Double Logarithmic Approximation (DLA) of the generating function evolution equation. The resulting modified DLA (MDLA) expressions deviate markedly from the DLA results, supporting earlier conclusions that the latter fail to provide a reliable description of experimental data Malaza and Webber (1984); Dokshitzer (1993); Duan et al. (2025); Dokshitzer and Webber (2025a, b), although recent DLA-based studies have successfully explained the observed KNO scaling in DIS Liu et al. (2023a, b). For gluon jets, our results are consistent with ref. Dokshitzer (1993), while for both quark and gluon jets they exhibit quantitative differences compared with the recently suggested QCD-inspired expressions in ref. Dokshitzer and Webber (2025b).
@@
The MDLA KNO scaling function for gluon jets depends solely on the anomalous dimension $\gamma_{0}$, while the quark KNO scaling function additionally depends on the ratio of mean multiplicities of quark and gluon jets in the asymptotic high-virtuality limit, $c_{q}$. Inspired by recent studies in refs. Dokshitzer and Webber (2025a, b), we treat $c_{q}$ as a fitting parameter in our analysis of jets in $pp$ collisions, rather than using the DLA value of $C_{F}/C_{A}$. We find that by fixing $\gamma_{0}$ and $c_{q}$, the MDLA results can very well reproduce the approximate KNO scaling functions for both quark and gluon jets in the two leading jets of $pp$ collisions at $\sqrt{s}=13$ TeV, as extracted using PYTHIA for a broad jet $p_{T}$ range of 0.10.12.52.5 TeV in ref. Duan et al. (2025). Furthermore, we observe that combining the MDLA scaling functions with the leading-order cross section and the N3LO mean multiplicities in refs. Dremin and Gary (1999); Capella et al. (2000) provides a very good description of the inclusive charged-particle multiplicity distributions measured by ATLAS Aad and others (2019).
@@
Describing the inclusive multiplicity distributions requires additional theoretical input, such as the fraction of jets initiated by each parton type and mean multiplicities. This extra dependence may introduce uncertainties when evaluating the validity of KNO scaling. To address this and as a consistency check, we utilize jet substructure techniques to distinguish quark and gluon jets. Specifically, we apply the topic modeling method Metodiev and Thaler (2018); Komiske et al. (2018) to two jet samples provided by ATLAS Aad and others (2019) to extract multiplicity distributions for jet topics, serving as an approximation for quark and gluon jets. We find that these results are consistent with our MDLA results, although the experimental uncertainties remain significant.
@@
Several aspects of our study could be further improved. First, a complete calculation of the KNO scaling functions for quark and gluon jets beyond the DLA approximation is still lacking. This is important because, although MDLA only incorporates energy conservation compared to DLA, it already introduces significant modifications. Moreover, describing the high-multiplicity tail observed by ATLAS Aad and others (2019), where hadronic multiplicities exceed the average by a factor of 3 or more, would require detailed knowledge of hard jet constituents Dokshitzer and Webber (2025a), which lies beyond both DLA and MDLA. A key issue to address beyond DLA is how to perform a direct calculation using the same jet definition (e.g., the anti-$k_{t}$ algorithm) employed in experiments. For inclusive multiplicity distributions, higher-order dijet cross sections are available Currie et al. (2017). In addition, the discrepancies between theoretical predictions and the measured mean multiplicities need to be resolved to achieve better agreement. One promising approach for high-precision calculations involves infrared-safe multiplicity definitions Medves et al. (2022, 2023), although multiplicity distributions within this approach are yet to be studied. Notably, KNO scaling functions are infrared safe within both DLA and MDLA, and parton-level results have been used for comparisons with experimental data in this work (see also refs. Dokshitzer and Webber (2025a, b)). At the same time, non-perturbative effects are observed in PYTHIA simulations Vertesi et al. (2021); Duan et al. (2025), highlighting the importance of clarifying the role of hadronization effects Azimov et al. (1985); Kang et al. (2024). Another promising avenue is to explore how studies of multiplicity distributions in quark and gluon jets can benefit from complementary methods of quark–gluon discrimination Gallicchio and Schwartz (2011, 2013); Larkoski et al. (2014); Gras et al. (2017); Metodiev et al. (2017); Dreyer et al. (2022). On the experimental side, more precise measurements are needed, and the $n_{ch}$ binning should be refined.
@@ Acknowledgements.
X.D. thanks Dr. Wen-Ya Wu for valuable discussions regarding the treatment of experimental data. This work is supported by the European Research Council under project ERC-2018-ADG-835105 YoctoLHC; by Maria de Maeztu excellence unit grant CEX2023-001318-M and projects PID2020-119632GB-I00 and PID2023-152762NB-I00 funded by MICIU/AEI/10.13039/501100011033; and by ERDF/EU. It has received funding from Xunta de Galicia (CIGUS Network of Research Centres). X.D. is supported by the China Scholarship Council under Grant No. 202306100165. L.C. is supported by the Marie Skłodowska-Curie Actions Postdoctoral Fellowships under Grant No. 101210595. X.D. and G.M. are supported by the National Natural Science Foundation of China under Grants No.12147101, No. 12547102, No. 12325507, the National Key Research and Development Program of China under Grant No. 2022YFA1604900, and the Guangdong Major Project of Basic and Applied Basic Research under Grant No. 2020B0301030008. B.W. acknowledges the support of the Ramón y Cajal program with the Grant No. RYC2021-032271-I and the support of Xunta de Galicia under the ED431F 2023/10 project.
@@ Appendix A Mean multiplicities at N3LO
To clarify the role of the parameters appearing in the N3LO results Dremin and Gary (1999); Capella et al. (2000), we provide a summary in this appendix. The mean multiplicity for gluon jets takes the form
@@
$r_{0}=\frac{N_{c}}{C_{F}},\qquad\delta_{q}(y)=\delta_{g}(y)+\frac{C}{\sqrt{y}}r_{1}+\frac{C^{2}}{y}(r_{2}+\frac{r_{1}^{2}}{2}).$ (52)
The coefficients $a_{i},r_{i}$ depend on the number of flavors $n_{f}$ and are tabulated in refs. Capella et al. (2000). In our analysis, these coefficients are evaluated using $n_{f}=5$, with $a_{1}=0.314,a_{2}=-0.301,a_{3}=0.112,r_{1}=0.198,r_{2}=0.510$, and $r_{3}=-0.041$.